OpenStudy (anonymous):

What is HX(1/sqrt(2) |0> + 1/sqrt(2) |1>)?

OpenStudy (anonymous):

Actually the question is in other figure I just make my own question so I can have the basic idea how to solve this question.

OpenStudy (anonymous):

\[HX\left(\frac{1}{\sqrt(2)}|0\rangle+\frac{1}{\sqrt(2)}|+\rangle\right)\]